OpenStudy (loser66):

Does there exists a vector space V and a bilinear form w on V \(\oplus\)V such that w is not identically zero, but w (x,x) =0 for every x in V? Please, help

OpenStudy (loser66):

@wio

OpenStudy (loser66):

V = {M_2 = span {\(\left[\begin{matrix}1&0\\0&0\end{matrix}\right],\left[\begin{matrix}0&1\\0&0\end{matrix}\right],\left[\begin{matrix}0&0\\1&0\end{matrix}\right],\left[\begin{matrix}0&0\\0&1\end{matrix}\right]\)}} w \(\in V\oplus V~~ and ~~~~w (x,x) = x_1y_2 + x_2y_1\) , then w (x,x) =0 for all x My question is : is there any other vector space else?

OpenStudy (anonymous):

maybe something like \(w(x,y) = x-y\)? But it has to be bilinear

OpenStudy (loser66):

yes, I think so, but (x, x,) not (x, y) I think of w (x,x) = x cross x , but not sure how to put it in the logic of bilinear form

OpenStudy (loser66):

Can I do : in \(R^2\) w (x,x) = x1y2 - x2 y1=0 for any x, let check. x =(x1, x2) so, w ( x, x ) = \(x_1x_2 - x_2x_1 \) =0 for all x does it work?

OpenStudy (loser66):

oh, I have to go. Thanks for being here. :)